sir, i am not understanding the algebric identities.

Grade:8

2 Answers

venky
8 years ago
In mathematics an identity is an equality relation A = B, such that A and B contain some variables and A and B produce the same value as each other regardless of what values (usually numbers) are substituted for the variables. In other words, A = B is an identity if A and Bdefine the same functions. This means that an identity is an equality between functions that are differently defined. For example (a + b)2 =  a2 + 2ab + b2 and cos2(x) + sin2(x) = 1 are identities. Identities are sometimes indicated by the triple bar symbol ≡ instead of =, theequals sign.
OUTPASS
8 years ago
This means that an identity is an equality between functions that are differently defined. For example (a + b)2 =  a2 + 2ab + b2 and cos2(x) + sin2(x) = 1 are identities.